OpenStudy (anonymous):

log3(5)+log3x=log3(10)

mathslover (mathslover):

ok so here base is same

mathslover (mathslover):

\[\large{log_a(b)+log_a(c)=log_a(bc)}\]

mathslover (mathslover):

can u write the given question in this form ?

OpenStudy (anonymous):

log3(5)+log3x=log3(10) => log3 (5 X 3x ) =log3(10) =>log3(15x) =log3(10) =>15x=10 =>x=2/3

mathslover (mathslover):

i think u did a mistake some where @DHASHNI \[\large{log_3(5)+log_3(x)=log_3(5*x)}\]but it is given equal to \(\large{log_3{10}}\) \(\large{\log_3(5x)=log_3(10)}\) \(\large{5x=10}\) \(\large{x=\frac{10}{5}=2}\)

OpenStudy (anonymous):

yes

mathslover (mathslover):

got it @cypress

OpenStudy (anonymous):